8. Cribb Corporation uses direct labor-hours in its predetermined overhead rate. At the beginning of the...

8.

Cribb Corporation uses direct labor-hours in its predetermined overhead rate. At the beginning of the year, the estimated direct labor-hours were 13,800 hours and the total estimated manufacturing overhead was $353,280. At the end of the year, actual direct labor-hours for the year were 13,500 hours and the actual manufacturing overhead for the year was $340,650. Overhead at the end of the year was: (Round your intermediate calculations to 2 decimal places.)
Answer #1

Predetermine overhead rate = Estimated overhead/estimated labor hour

= 353280/13800

Overhead rate = 25.60 per hour

Overhead applied = 13500*25.60 = 345600

Actual overhead = 340650

Over applied overhead at the end of the year = 345600-340650 = 4950

Overhead at the end of the year was : Overapplied by $4950